### v1

# FedReFT: Federated Representation Fine-Tuning with All-But-Me Aggregation

Abstract
[…]
FedReFT: Federated Representation Fine-Tuning with All-But-Me Aggregation
Fatema Siddika1thanks: Equal contribution, Md Anwar Hossen [^1] footnotemark: 1, J. Pablo Muñoz2, Tanya Roosta3, Anuj Sharma1, Ali Jannesari1 1Iowa State University 2Intel Labs, 3University of California, Berkeley {fatemask, manwar, anujs, jannesar}@iastate.edu jpablomch@gmail.com, tanya.roosta@gmail.com
1 Introduction
Fine-tuning has emerged as a core strategy for adapting large language models (LLMs) to various downstream tasks, allowing for a broad generalization from minimal task-specific data Ding et al. (2023); Ziegler et al. (2019). However, traditional fine-tuning is computationally expensive and memory-intensive, which poses scalability challenges. This is further amplified in resource-constrained environments, such as smartphones, where full model updates are often infeasible due to limited resources. To address these challenges, parameter-efficient fine-tuning (PEFT) methods such as Adapter Tuning Houlsby et al. (2019), BitFit Zaken et al. (2022), Prefix Tuning Li and Liang (2021), Prompt Tuning Lester et al. (2021), and Low-Rank Adaptation (LoRA) Hu et al. (2021a), have been proposed, significantly reducing the cost of adaptation by updating only a small subset of model weights.
PEFT has emerged as the preferred method for efficiently adapting large language models (LLMs) without sacrificing performance. However, most PEFT approaches assume centralized data access, which is unrealistic in many real-world scenarios where data is distributed across users or devices with varying tasks and privacy concerns. Federated Learning (FL) offers a solution by enabling collaborative model training without centralizing data, but prior FL work often emphasizes task-specific tuning rather than learning generalizable representations. In practice, clients frequently work on diverse or specialized tasks, making global representation learning both more difficult and more essential.
While PEFT typically modifies model weights, recent interpretability research highlights the potential of hidden representations, which encode rich semantic information. Representation Fine-Tuning (ReFT) Wu et al. (2024b) leverages this by directly intervening in hidden layers, achieving stronger performance than methods like LoRA. Despite ReFT’s success in centralized settings, it has yet to be adapted for FL, where challenges such as data heterogeneity, varying model capacities, and limited computational resources complicate aggregation and reduce effectiveness. In order to study the challenges of representation-level fine-tuning under heterogeneous federated settings, and to evaluate the effectiveness of our proposed aggregation strategy, we put forward the following research questions:
1) How can we aggregate representation-level updates in Federated Learning without compromising semantic alignment across task-heterogeneous clients?
2) Is weighted averaging sufficient for aligning semantically rich hidden representations, or is a more robust and personalized strategy needed to preserve local semantics while leveraging global knowledge?
To address these questions, we introduce Federated Representation Fine-Tuning (FedReFT), a novel framework for personalized and parameter-efficient federated representation fine-tuning. FedReFT builds on the core idea of ReFT by enabling clients to inject lightweight intervention components (i.e., sparse low-rank matrices $W$, $R$, $b$) directly into hidden representations, making it particularly useful for edge devices with limited resources. To mitigate the degradation in semantic alignment caused by naive aggregation such as vanilla weighted average aggregation (FedAvg) McMahan et al. (2017), we propose the All-But-Me (ABM) aggregation strategy. Instead of averaging all client updates uniformly, ABM constructs a personalized global intervention for each client by computing the geometric median over updates from all other clients.
The key contributions of our work are as follows: Contribution 1: We address a critical gap in the utilization of ReFT in FL setting by introducing a novel aggregation strategy, All-But-Me (ABM), specifically designed for low-rank, representation-level interventions. ABM addresses the challenge of semantic misalignment caused by naive aggregation strategies, preserving client-specific semantics while enabling stable collaboration.
Contribution 2: We propose FedReFT, a novel framework for personalized and parameter-efficient federated fine-tuning based on representation-level interventions. Contribution 3: We evaluate the framework by simulating task heterogeneity, i.e., assigning different tasks to clients, all derived from a common dataset. This setup mimics real-world scenarios where clients pursue distinct objectives over structurally similar data, allowing us to evaluate the effectiveness of FedReFT and ABM under realistic conditions. Paper outline: The remainder of the paper is organized as follows. Section 2 formally defines the problem setting, the motivation behind our work and the challenges posed by heterogeneous FL setting. Section 3 details our methodology, including the FedReFT mechanism and the All-But-Me (ABM) aggregation strategy. Section 4 presents our experimental results and evaluates the effectiveness of our approach on multiple benchmarks. Section 5 concludes the paper with key insights and future directions. We defer additional details to the appendices.
Figure 1: Illustration of the relationship between the average accuracy and trainable parameter (in $\%$)for various federated PEFT methods on Commonsense, Arithmetic, and GLUE benchmarks using LLaMA-3.2B, LLaMA-3 8B, and RoBERTa-large models, respectively. FedReFT achieves competitive or state-of-the-art performance while training significantly fewer parameters, resulting in improved communication efficiency and reduced transmission cost in federated learning settings.
2 Problem Formulation and Motivation
In this section, we present the motivation to apply ReFT in FL and formalize the problem of adapting personalized representation. Although ReFT offers parameter-efficient updates in the representation space, its application in FL faces key challenges, including task heterogeneity, semantic misalignment, and unstable aggregation among heterogeneous clients. We highlight these challenges and formulate the objective of enabling parameter-efficient and semantically aligned adaptation in FL using ReFT. Challenge 1: LoReFT in FL Settings: ReFT  Wu et al. (2024b) offers an attractive alternative by modifying hidden activations instead of model weights. By intervening directly in structured semantic subspaces, ReFT supports interpretable, modular, and task-aligned adaptation, particularly advantageous in task-heterogeneous FL settings. However, full ReFT incurs considerable communication costs and poses integration challenges when clients use different model capacities or architectures.
To bridge this gap, we adopt Low-Rank Linear Subspace ReFT (LoReFT)Wu et al. (2024b), a lightweight ReFT variant that constrains interventions to a learnable low-rank subspace. This design significantly reduces overhead while maintaining semantic control, making it a promising candidate for FL. We follow the LoReFT intervention formulation from Wu et al. (2024b) on hidden representations $h\in\mathbb{R}^{d}$ which is defined as:
$\Phi_{\text{LoReFT}}(h)=h+R^{\top}(Wh+b-Rh),$ (1)
where, $W\in\mathbb{R}^{r\times d}$ is a low-rank projection matrix with $d$ as the representation dimension and $r$ as the subspace intervention dimension, $R\in\mathbb{R}^{r\times d}$ is a low-rank projection matrix with orthonormal rows, and $b\in\mathbb{R}^{r}$, with $r\ll d$. This structure, inspired by Distributed Interchange Intervention (DII) Geiger et al. (2024), enables semantically grounded, low-rank adaptation suitable for scalable and privacy-preserving FL. Despite its efficiency, applying LoReFT in FL raises several non-trivial challenges: LoReFT modifies internal representations that are sensitive to client-specific data distributions. Aggregating these interventions naïvely using FedAvg can cause semantic interference or collapse. Without global synchronization, low-rank updates may evolve in divergent directions, especially when tasks are dissimilar. Applying shared LoReFT interventions across clients risks overfitting to shared patterns while ignoring local semantics. Considering all these challenges, the major research question is:
Can representation-level adaptation via LoReFT achieve personalization and stability in federated environments without collapsing under task and data heterogeneity?
FedReFT uses All-But-Me(ABM) aggregation to robustly combine intervention parameters while preserving personalization in heterogeneous FL. Challenge 2: Federated Fine-Tuning under Task Heterogeneity: A central motivation of our work is to address task heterogeneity in real-world FL, where clients perform fundamentally different tasks rather than optimizing a shared objective. For example, clients may work on distinct reasoning tasks within natural language QA that demand different semantic skills. While centralized fine-tuning has proven effective for such tasks, it assumes access to all data, which is unrealistic in decentralized settings. In FL, each client sees only a local, task-specific subset of the broader reasoning space, leading to highly heterogeneous training distributions, a common challenge in multi-department or cross-domain deployments. This raises the question:
How can we learn a global representation that generalizes across tasks when each client trains only on a fragment of the broader task distribution?
Standard methods like FedAvg McMahan et al. (2017) struggle in this regime, as they average semantically misaligned updates, often resulting in degraded performance or collapsed representations. Formally, let each client $i$ have a dataset $\mathcal{T}_{i}=\{X_{i},Y_{i}\}$ and optimize a personalized model $\boldsymbol{\theta}_{i}$ by solving:
$\min_{\boldsymbol{\Theta}}\ \frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(X_{i},Y_{i},\boldsymbol{\theta}_{i}),$ (2)
where $\mathcal{L}$ is the task-specific loss and $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{i}\}_{i=1}^{N}$ is the set of client-specific models. Our proposed method, FedReFT, can successfully address this research challenge. FedReFT enables scalable, personalized representation learning across heterogeneous tasks, allowing global reasoning capabilities to emerge from decentralized, task-specific updates.
Table 1: Comparison of different aggregation methods of GLUE task on ROBERTa, Commonsense and Arithmetic reasoning task on LLaMA-2 7B.
Task FedAvg (%) FedReFT (%)
Commonsense 70.16 70.77
Arithmetic 15.36 17.21
GLUE 88.17 89.77
Challenge 3: Learnable Parameter Sharing with the Server: When applying ReFT Wu et al. (2024b) in a FL setting from the perspective of learnable parameter sharing, a fundamental question is:
Which of these parameters should be communicated to the server for collaborative aggregation?
Figure 2: FedReFT with ABM Aggregation. Clients cross-task demonstrate personalization while maintaining alignment with the global representation. (1)-(2): Each client applies LoReFTWu et al. (2024b) interventions to train learnable parameter {$W$, $R$, bb} to modify hidden representations $h$ in a low-rank edit subspace. (3): Clients fine-tune {$W$, $R$, bb} locally and partially fuse received All-But-Me aggregated updates with their own. (4): The server performs ABM aggregation using the geometric median over other clients’ intervention parameters to generate $W_{k}^{\text{ABM}},R_{k}^{\text{ABM}},b_{k}^{\text{ABM}}$.
In FedReFT, each client fine-tunes hidden representations by introducing learnable low-rank intervention parameters $W$, $R$, and a bias $b$ into a frozen backbone model. Sharing only some part of the intervention parameters leads to incomplete information transfer and breaks the low-rank structure critical for generalization. $W$ projects representations into a low-dimensional space, and $R$ reconstructs them; omitting either disrupts compositionality and limits alignment across heterogeneous clients. In Table 2, empirical results show that partial sharing significantly degrades performance and representation alignment under task heterogeneity. Therefore, FedReFT shares the full set of learnable intervention parameters $(W,R,B)$ from each client with the server.
3 Methodology
In this section, we introduce FedReF+, designed to address the challenges we discussed in the previous section. An illustrative overview of FedReFT is shown in Figure 2.
3.1 Intervention Parameter Sharing Strategies
To reduce communication overhead while maintaining personalization, we explore three strategies for sharing local intervention parameters with the server. These strategies are summarized in Table 2 and represent different trade-offs between expressiveness and communication efficiency:
1) Full Intervention Sharing: $\mathbf{W}\in\mathbb{R}^{r\times d}$, $\mathbf{R}\in\mathbb{R}^{r\times d}$, $\mathbf{b}\in\mathbb{R}^{r}$} This strategy shares the complete set of intervention parameters, capturing client-specific compression ($\mathbf{W}$), transformation ($\mathbf{R}$), and translation ($\mathbf{b}$). It enables the most accurate reconstruction of local updates and yields the best global performance, especially under high heterogeneity.
2) No Bias Sharing: $\mathbf{W}\in\mathbb{R}^{r\times d}$, $\mathbf{R}\in\mathbb{R}^{r\times d}$} This variant omits the bias term $\mathbf{b}$ but retains the directional transformation via $\mathbf{W}$ and $\mathbf{R}$. While it allows the server to align low-rank subspace transformations across clients, it lacks the ability to model per-dimension translation shifts, which can hinder fine-grained personalization. 3) No $\mathbf{W}$ Sharing: $\mathbf{R}\in\mathbb{R}^{r\times d}$, $\mathbf{b}\in\mathbb{R}^{r}$} This configuration excludes $\mathbf{W}$, giving the server access only to the reconstruction and shift parameters. Without knowledge of how the local signals were encoded, the server’s ability to interpret or align updates is severely limited. The $\{\mathbf{W},\mathbf{R},\mathbf{b}\}$ strategy provides the highest fidelity for aggregation, $\{\mathbf{W},\mathbf{R}\}$ offers a balanced compromise, and $\{\mathbf{R},\mathbf{b}\}$ prioritizes communication efficiency at the cost of semantic alignment and global performance.
Table 2: Performance vs. parameter efficiency for different LoReFT sharing strategies (Uplink) for $C$ clients on commonsense reasoning task following the second experiment design. GLUE task on ROBERTa, Arithmetic and Commonsense on LLaMa-2 7B model.
Task Strategy TP(% $\downarrow$) Score
GLUE W,R,b 0.01384 94.31
W,R 0.01383 64.03
R,b 0.00693 74.12
Arithmetic W,R,b 0.03114 29.01
W,R 0.03114 26.13
R,b 0.01557 25.77
Commonsense W,R,b 0.03114 73.82
W,R 0.03114 70.63
R,b 0.01557 68.02
3.2 Intervention Design for Federated Classification Tasks
Following the formulation in ReFT Wu et al. (2024b), for a given client, we define the classification head $H_{\psi}$ with parameters $\psi=\{W_{o},b_{o},W_{d},b_{d}\}$ operates on the CLS token representation $z\in\mathbb{R}^{d}$ from the final layer:
$H_{\psi}(z)=\text{softmax}\left(W_{o}\cdot\tanh(W_{d}z+b_{d})+b_{o}\right).$ (3)
We jointly optimize the intervention parameters $\phi$ and the classifier $\psi$ using cross-entropy loss over input $x$ and label $y$:
$\min_{\phi,\psi}\left\{-\log H_{\psi}\left(y\mid z_{\phi}(x)\right)\right\}.$ (4)
3.3 All-But-Me (ABM) Aggregation
In heterogeneous FL, the integration of shared knowledge without compromising local task-specific adaptation remains a core challenge. Standard aggregation methods such as FedAvg McMahan et al. (2017), which averages client models into a single global model, are often suboptimal in non i.i.d. scenarios. They risk overwriting valuable client-specific representations and rely on fixed mixing weights that may further reduce personalization. Table 1 depicts the comparison. To overcome these limitations, we propose the All-But-Me (ABM) aggregation strategy. Instead of initializing clients with a global model, each client continues to update its local parameters while partially incorporating knowledge aggregated from other clients. Specifically, each client $k$ receives a robustly aggregated set of intervention parameters $\{\mathbf{W}^{\text{ABM}}_{k},\mathbf{R}^{\text{ABM}}_{k},\mathbf{B}^{\text{ABM}}_{k}\}$, calculated from the updates of all other clients using a geometric median:
$\mathbf{W}_{k}^{\text{ABM}}=\text{ABM}\left(\left\{\mathbf{W}_{m}^{\text{local}}\right\}_{m\neq k}\right),$ (5)
$\mathbf{R}_{k}^{\text{ABM}}=\text{ABM}\left(\left\{\mathbf{R}_{m}^{\text{local}}\right\}_{m\neq k}\right),$
$\mathbf{B}_{k}^{\text{ABM}}=\text{ABM}\left(\left\{\mathbf{B}_{m}^{\text{local}}\right\}_{m\neq k}\right).$
The client then performs a personalized update by interpolating between its local parameters and the ABM-aggregated ones using a mixing factor $\alpha\in[0,1]$. We have discussed in detail how the evaluation loss-based $\alpha$ tuning works in the appendix D.
$\mathbf{X}_{k}^{\text{new}}$ $=(1-\alpha)\cdot\mathbf{X}_{k}^{\text{local}}+\alpha\cdot\mathbf{X}_{k}^{\text{ABM}},$ (6)
$\mathbf{X}$ $\in\{\mathbf{W},\mathbf{R},\mathbf{B}\}.$
Before the next local training using the $\mathbf{R}_{k}^{\text{new}}$, we do the orthogonal transformation of $\mathbf{R}_{k}^{\text{new}}$ to keep the original property of $R$.
ABM via Geometric Median.
The geometric median (also known as the spatial or $L_{1}$ median) offers a robust alternative to the arithmetic mean, particularly under client heterogeneity and adversarial conditions Maronna and Martin (2006); Weiszfeld (1937). Given a set of vectors $\mathcal{S}=\{x_{1},\ldots,x_{n}\}\subset\mathbb{R}^{d}$, it is defined as:
$x^{*}=\arg\min_{x\in\mathbb{R}^{d}}\sum_{i=1}^{n}\|x-x_{i}\|_{2},$ (7)
which minimizes the sum of Euclidean distances to all elements in the set. This estimator is robust to outliers and misaligned updates, making it well-suited for federated settings. We instantiate the ABM function using the geometric median, where each client $k$ receives an aggregated intervention vector computed from $\mathcal{S}_{k}=\{x_{m}\}_{m\neq k}$:
$\text{ABM}(\mathcal{S}_{k})=\arg\min_{x\in\mathbb{R}^{d}}\sum_{x_{m}\in\mathcal{S}_{k}}\|x-x_{m}\|_{2}.$ (8)
To solve this optimization efficiently, we employ Weiszfeld’s algorithm Weiszfeld (1937), an iterative method known to converge under mild conditions. Details of the algorithm are provided in Appendix E. By avoiding direct averaging and incorporating semantically meaningful low-rank intervention updates through robust aggregation, ABM enables each client to benefit from the knowledge of others without sacrificing local personalization. This approach enhances stability and generalization across non-i.i.d. and task-heterogeneous FL environments.
Table 3: Federated fine-tuning performance of LlaMa-3.2 3B across five commonsense reasoning tasks with MT experimental setup where clients train on heterogeneous task mixtures to promote generalizable representations. *Performance results of all baseline methods and the experimental setup are taken from Singhal et al. (2025).
Method R # TP(M) $\downarrow$ TP(%) BoolQ PIQA SIQA HellaS. WinoG Avg $\uparrow$
FLoRA* 32 243.15 7.58 65.05 82.81 74.67 81.84 76.01 78.83
FedIT* 32 48.63 1.51 62.99 81.50 73.13 76.83 71.51 75.74
FFA-LoRA* 32 24.31 0.76 62.87 80.03 68.53 70.02 65.56 71.11
Fed-SB * 120 2.83 0.0884 64.86 81.66 74.87 81.67 75.22 75.66
FedReFT (Ours) 4 1.38 0.0428 63.09 82.10 72.36 90.27 69.22 75.41
8 2.75 0.0857 64.01 81.18 72.11 90.71 71.01 75.66
16 6.194 0.1927 63.42 81.61 73.64 91.23 71.35 76.05
32 11.01 0.3427 64.53 81.34 73.39 91.51 71.32 76.22
FedReFT (tie $\phi$, Ours) 4 0.688 0.0214 49.94 81.23 72.72 89.84 68.43 72.43
8 1.38 0.0428 57.15 81.22 72.77 90.56 68.50 74.04
Table 4: Performance comparison across arithmetic reasoning tasks with the Distict Task and Mixed Task setup.
FedReFT Distinct Task (DT) Mixed Task (MT)
Models AQuA GSM8K SVAMP Avg $\uparrow$ AQuA GSM8K SVAMP Avg $\uparrow$
LLaMa 7B 25.59 25.47 49.80 33.62 22.83 14.33 27.10 21.42
LLaMa-2 7B 29.53 32.45 57.3 39.76 21.65 20.39 31.5 24.51
LLaMa-3 8B 34.64 48.98 73.60 52.41 31.89 48.90 70.04 50.48
4 Experimental Validation
To evaluate FedReFT, we conduct extensive experiments on three different NLP benchmarks covering over 12 datasets. Our objective is to present a comprehensive assessment of how this approach performs in various NLP tasks. We experiment with both masked and autoregressive language models, including RoBERTa-large Liu et al. (2019), TinyLlama-1B Community (2023), LLaMA 7B Touvron et al. (2023a), LLaMA-2 7B and 13B Touvron et al. (2023b), LLaMA-3.2B and LLaMA-3 8B AI (2024), across multiple settings and scales. Our comparisons include state-of-the-art baselines, such as LoRAHu et al. (2021b), FedIT Zhang et al. (2024), FFA-LoRASun et al. (2024), FedDPA-LoRALong et al. (2024), FedSA-LoRAGuo et al. (2024), Fed-SB Singhal et al. (2025) and FLoRAWang et al. (2024) focusing on both parameter efficiency and performance trade-offs. We align the experimental setup configurations with the baseline papers to ensure fair comparisons. To optimize memory usage, we load all base language models with torch.bfloat16 precision. All experiments are executed on a single NVIDIA A100-SXM4-80GB GPU, except for LLaMA-2 13B, which is run on a GPUH200x8 141GB system to accommodate the computational demands of large-scale federated fine-tuning. The results are averaged over two runs to report the mean performance.
Hyperparameter Configuration. In the experiments, we determine how many interventions to learn, as well as which input positions and layers to apply them to. We apply interventions at a fixed number of layers $L$, and at prefix ($p$) and suffix ($s$) positions in the input prompt. We narrow the hyperparameter search space for Federated Learning by adopting the configuration used in the centralized ReFTWu et al. (2024b) paper. The appendix B provides a brief overview of the hyperparameter search space. We experiment with whether to share (tie) the intervention parameters $\phi$ across different input positions within the same layer. Given the positions $P=\{1,\ldots,p\}\cup\{n-s+1,\ldots,n\}$, we define the untied and tied variants Wu et al. (2024b):
$\mathbf{I}_{\text{untied}}=\left\{\langle\Phi,\{p\},l\rangle\mid p\in P,\,l\in L\right\},$ (9)
$\qquad\mathbf{I}_{\text{tied}}=\left\{\langle\Phi,P,l\rangle\mid l\in L\right\}.$
Task Distribution Rationale.
We design two experimental setups to study how global representations converge under diverse task distributions. In the Mixed-Task (MT) setup, each client trains on a subset of a combined reasoning dataset but is evaluated on a single task, encouraging generalized, transferable representations through ABM aggregation. This reflects collaborative learning across varied yet related tasks. In the Distinct Task (DT) setup, each client trains on a unique reasoning task, enabling personalized fine-tuning while still leveraging global updates. Despite higher task heterogeneity, this setup maintains stable performance as model capacity increases. Both setups show that FedReFT supports effective generalization in MT and robustness in DT.
4.1 Commonsense Reasoning
We evaluate global representation generation on eight commonsense reasoning tasks using the Commonsense170K dataset inspired by Singhal et al. (2025); Wu et al. (2024b). We use the same hyperparameter of  Singhal et al. (2025) and tune the intervention parameter in the Appendix B.1. This helps us tune important hyperparameters efficiently and also test their robustness across multiple commonsense reasoning tasks.
Datasets. For the first setup (MT design), we split the combined COMMONSENSE170K Hu et al. (2023) commonsense reasoning tasks among clients and use them for fine-tuning. Each client evaluates one of the commonsense reasoning tasks. BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), and WinoGrande Sakaguchi et al. (2021). For the second setup (DT design), each client fine-tunes on only one of these five commonsense reasoning tasks and is evaluated using the same task. All examples are formatted as multiple-choice questions, requiring the model to directly generate the correct answer without providing rationales. We adopt the prompt template from Hu et al. Hu et al. (2023) with minor modifications, including additional string normalization by removing leading and trailing whitespace.
Table 5: Performance comparison across GLUE Tasks on RoBERTa model for $C=3$, FedReFT use rank rank 1. *Performance results of all baseline methods are taken from Guo et al. (2024) and use LoRA rank 8.
Setup Method # TP(M) $\downarrow$ TP(%) MNLI-m SST-2 QNLI QQP Avg $\uparrow$
Standalone FT 355 100 88.8 96.0 93.8 91.5 91.87
LoRA* 1.83 0.515 88.71 95.16 91.16 85.33 89.33
LoReFT 0.053 0.015 89.2 96.2 94.1 88.5 92.0
FL FFA-LoRA* 1.44 0.405 88.83 94.95 91.52 86.71 89.39
FedDPA-LoRA* 2.62 0.737 88.99 95.50 90.74 85.73 89.47
FedSA-LoRA* 1.83 0.551 90.18 96.00 92.13 87.48 90.43
FedReFT (ours) 0.053 0.015 88.86 95.17 94.52 86.57 90.93
[…]
Table 6: Performance comparison on arithmetic reasoning tasks for GSM8K on LLaMa-3 8B model with LoRA rank 8, where clients enable consistent evaluation of representation generalization. *Performance results of all baseline methods are taken from Guo et al. (2024).
Method # TP(M) $\downarrow$ TP(%) GSM8K
LoReFT 4.19 0.052 48.33
LoRA* 30.40 0.38 46.23
FedSA-LoRA* 30.40 0.38 46.63
FFA-LoRA* 15.2 0.19 46.32
FedReFT (tie $\phi$) 2.09 0.0261 49.35
FedReFT 4.19 0.0622 49.68
4.2 Arithmetic Reasoning
For the arithmetic reasoning tasks, we design three experimental settings to fine-tune models on various arithmetic reasoning tasks. We follow the same hyperparameter tuning strategy as used in COMMONSENSE170K in Appendix B.1, which uses a development set to select the best-performing configuration. Evaluation is based solely on the final numeric or multiple-choice answer, disregarding intermediate reasoning steps.
Datasets. In the first setting following MT, we split a combined arithmetic reasoning dataset, MATH10K Hu et al. (2023) which includes four arithmetic reasoning tasks with chain-of-thought solutions generated by a language model. Each client reports performance using test set one of three tasks: AQuA Ling et al. (2017), GSM8K Cobbe et al. (2021), and SVAMP Patel et al. (2021). In the second setting following DT, each client is assigned one arithmetic reasoning task for both fine-tuning and evaluation. Both MT and DT setup results are reported in Table 4. In the third setting following Guo et al. (2024); Kuang et al. (2024), we split the dataset GSM8K into three clients under an IID distribution, and the results are shown in Table LABEL:tbl:math_setup_3_baseline. All optimization hyperparameters remain consistent with that setup.
Results. In Table LABEL:tbl:math_setup_3_baseline, FedReFT demonstrates strong performance while using significantly fewer trainable parameters than existing baselines. Notably, FedReFT achieves the highest accuracy among all methods, while FedReFT (tie $\phi$) offers a compelling trade-off between performance and efficiency. These results highlight the scalability and efficiency of our representation-tuning approach. The Distinct Task (DT) setup represents task heterogeneity. Hence, fine-tuning allows clients to learn highly personalized, task-specific representations while benefiting from global aggregation. As a result, the DT setup yields higher performance, as seen in Table 4. In contrast, the MT setup, where clients train on heterogeneous task mixtures to promote global generalizable representations. This blending of tasks during fine-tuning leads to general representation learning but can degrade performance on specific evaluation tasks due to misaligned representation and conflicting task objectives.
4.3 Natural Language Understanding
We evaluate the effectiveness of FedReFT in learning generalizable representations for Natural Language Understanding (NLU) using the GLUE benchmark Wang et al. (2018). The objective is to fine-tune NLU to learn global representations that capture task-level semantics. By aligning intermediate representations for downstream classification performance. This setup allows us to test whether lightweight intervention tuning can align representations across clients within a single NLU task.
Hyperparameter Tuning. We tune hyperparameters separately for each task, following common practice for PEFT methods Hu et al. (2023) in FL. To reduce the impact of random seed variability, we perform hyperparameter tuning using a fixed seed and report the average performance across that seed and two additional unseen seeds. Details are provided in the Appendix B.2.
Results. Table LABEL:tbl:glue_task depicts that our approach performs strongly across GLUE tasks while using very few trainable parameters. It performs competitive or outperforms other methods, showing that it can learn good representations even in a federated setting. Despite using over 30× fewer parameters than some baselines, it still achieves competitive results, making it both efficient and effective.
4.4 Ablation Studies
We conduct ablation studies to better understand the effectiveness of FedReFT, specifically examining the role of geometric median-based All-But-Me aggregation. Details of this analysis are provided in Appendix F.
5 Conclusion
In this work, we addressed a critical gap in the deployment of Representation Fine-Tuning within Federated Learning settings by proposing a novel aggregation strategy tailored to its low-rank, representation-level interventions. While ReFT improves upon traditional PEFT methods like LoRA by operating on semantically rich hidden representations, its application in FL is limited by data heterogeneity, model diversity, and the shortcomings of standard aggregation methods. To address these, we propose All-But-Me aggregation, enabling clients to adapt their local ReFT parameters using a robust average of others’ interventions. FedReFT ensures parameter efficiency and semantic alignment. Extensive experiments under task heterogeneity and different heterogeneous settings show that ABM consistently enhances convergence, generalization, and robustness, making it a practical and effective solution for personalized representation learning in federated systems.
Limitations
Due to computational constraints, our current study focuses primarily on LoReFT-based interventions within language models under a fixed set of hyperparameters. In future work, we aim to automate the parameter search space using a multi-agent coordination framework to better explore optimal low-rank configurations for each client. Although our current set-up does not explicitly address privacy, we are actively investigating how to integrate differential privacy mechanisms, such as DP-SGD, into the FedReFT framework without sacrificing personalization. Initial experiments in this direction are ongoing. Additionally, we are exploring the theoretical properties of ABM aggregation under adversarial or noisy clients, and whether it can be extended to other modalities beyond language, such as vision-language models in federated systems.
Appendix A Related Works
A.1 Parameter-Efficient Fine-Tuning (PEFT)
Fine-tuning LLMs is resource-intensive due to their large parameter counts. Parameter-efficient fine-tuning (PEFT) methods mitigate this by updating only a small subset of parameters while keeping pre-trained weights frozen Li and Liang (2021); He et al. (2021); Wang et al. (2022). Several PEFT approaches have been proposed, Adapter Tuning Houlsby et al. (2019), BitFit Zaken et al. (2022), Prefix Tuning Li and Liang (2021), Prompt Tuning Lester et al. (2021), and Low-Rank Adaptation (LoRA) Hu et al. (2021a). Among them, LoRA is widely adopted for its efficiency in approximating weight updates via low-rank matrices. Extensions such as ReLoRA Lialin et al. (2023) and RankAdapter Zhou et al. (2024) improve memory use and adapt ranks dynamically, though they lack theoretical guarantees. AdaZeta Yang et al. (2024) introduces zeroth-order optimization with convergence guarantees, while others Gao et al. (2024); Rajabzadeh et al. (2024); Valipour et al. (2022) explore adaptive ranks without formal proofs. LoRA has been integrated with Mixture-of-Experts models Li et al. (2024a); Wu et al. (2024a), as in AdaMoLE Liu and Luo (2024), to enable dynamic expert selection, and also with Neural Architecture Search for LLM compression Muñoz et al. (2025). These approaches primarily target weight updates, overlooking direct interventions in hidden representations, which are discussed next.
A.2 Representation Fine-Tuning (ReFT)
ReFT shifts fine-tuning from model weights to hidden representations, leveraging their semantic structure for efficient adaptation Wu et al. (2024b). Inspired by activation steering and representation engineering Avitan et al. (2024); Li et al. (2024b); Liu et al. (2023); Singh et al. (2024), ReFT enables task-specific control through fixed or learned interventions without updating the full model. Notably, Inference-Time Intervention (ITI) Li et al. (2024b) improves LLM truthfulness by modifying activations, while representation engineering Zou et al. (2023) combines representation reading and control for interpretable model behavior. Minimally Modified Counterfactuals (MMC) Singh et al. (2024) unify erasure and steering to reduce bias, and can be mapped to natural language edits Avitan et al. (2024), enhancing interpretability. These findings support direct representation manipulation as a lightweight and effective alternative to weight-based PEFT methods like LoRA.
A.3 Federated Fine-Tuning
Federated Learning (FL) McMahan et al. (2017) poses challenges for fine-tuning LLMs, including data heterogeneity, communication constraints, and model diversity. PEFT methods have emerged to address these issues efficiently Sun et al. (2022); Chen et al. (2022); Zhang et al. (2023). LoRA-based approaches such as FedLoRA Yi et al. (2023), Hyper-FloRA Lu et al. (2024), and Efficient FL Adapter Cai et al. (2023) offer modular and personalized adaptation across clients. Recent advances further incorporate privacy (FFA-LoRA Sun et al. (2024)), heterogeneous adaptation (FloRA Wang et al. (2024)), instruction tuning (FedIT Zhang et al. (2024)), and expert routing (DualFed Long et al. (2024), Sparse-FedMoE Tran et al. (2025)). In contrast, our proposed FedReFT shifts from weight updates to direct representation-level tuning via sparse intervention layers and introduces an All-But-Me (ABM) aggregation strategy to preserve semantic alignment while enabling robust knowledge sharing across non-IID clients.
A.4 Aggregation Methods in FL
To address the inherent heterogeneity and robustness challenges in federated learning, median-based aggregation strategies have been extensively studied as alternatives to simple averaging. Unlike the arithmetic mean, the geometric and coordinate-wise medians are significantly more resilient to outliers and adversarial updates, making them suitable for secure and personalized FL scenarios. For instance, coordinate-wise median aggregation has been proposed to defend against Byzantine clients in distributed optimization Blanchard et al. (2017). This was extended with geometric median-based gradient descent to improve statistical guarantees across diverse loss landscapes Yin et al. (2018). Further work demonstrated that coordinate-wise median and trimmed-mean-based methods achieve order-optimal convergence not only for strongly convex losses but also under non-strongly convex and even non-convex population losses Chen et al. (2017). Additionally, a one-round median-based algorithm was shown to maintain statistical optimality under quadratic convexity, offering a communication-efficient solution Chen et al. (2017). RFA Pillutla et al. (2022) maintains privacy and demonstrates improved robustness over standard averaging techniques, particularly in environments with high levels of data corruption. FedEAT Pang et al. (2025) integrates adversarial training in the embedding space with geometric median-based aggregation to enhance robustness while preserving performance. This work demonstrates that LoRA-based FL systems can effectively leverage geometric median aggregation. Inspired by these findings, we adopt geometric median aggregation in our FL framework to aggregate the All-But-Me (ABM) intervention parameter, weight $\mathbf{W}$, rotation $\mathbf{R}$, and bias $\mathbf{b}$. This provides stability across diverse client behaviors and loss geometries, improving personalization performance under data and objective heterogeneity.
Appendix B Hyperparameter Search Space
B.1 Hyperparameter Search Space for Commonsense and Arithmetic Reasoning
Following the ReFT framework Wu et al. (2024b), we construct a development set using the GSM8K dataset and consider only the last 300 samples. We trained the clients using LLaMa 7B model with the remaining training data and determined the best-performing hyperparameters based on the model’s performance on the development set. We further use this hyperparameter in another model directly. We set the maximum input sequence length to 512 tokens during training and tuning, and limit inference to 32 generated tokens. We use the same setup for commonsense reasoning with COMMONSENSE170k dataset. The hyperparameter search space is summarized in Tables LABEL:tab:appendix_gsm8k_hypertune and LABEL:tab:appendix_commonsense_hypertune.
During inference, we use greedy decoding (without sampling) for the commonsense reasoning benchmark, as it is a multi-token classification task. For arithmetic reasoning, we follow the decoding setup from  Hu et al. (2023), using a higher temperature of 0.3. This change helps avoid errors in HuggingFace’s decoding caused by unstable probabilities
Table 7: Narrow down the hyperparameter(HP) search space of LLaMA 7B models with FedReFT on the GSM8K development set, inspired from Wu et al. (2024b). The best-performing settings are underlined. We apply greedy decoding without sampling during hyperparameter tuning.
HP FedReFT
prefix+suffix, $p+s$ {p5+s5, p7+s7, p9+s9}
Tied weight $\phi$ {True, False}
Rank $r$ {8, 16, 32, 64}
Layer $L$ {all}
Dropout {0.00, 0.05}
Optimizer AdamW
[…]
LR scheduler Linear
Batch size {16, 32}
Warmup ratio {0.06, 0.10}
Clients {3, 5}
Epochs {3, 4, 5, 6}
Rounds 10
Table 8: Narrow down the hyperparameter (HP) search space of LLaMA 7B models with FedReFT on the COMMONSENSE170k development set, following the Appendix B.1. The best-performing settings are underlined. We apply greedy decoding without sampling during hyperparameter tuning.
HP FedReFT
prefix+suffix, $p+s$ {p5+s5, p7+s7}
Tied weight $p,s$ {True, False}
Rank $r$ {8, 16, 32, 64}
Layer $L$ {all}
Dropout {0.00, 0.05}
Optimizer AdamW
[…]
Weight decay {0}
LR scheduler Linear
Batch size {16, 32}
Warmup ratio {0.1}
Clients {3, 5}
Epochs {2, 3, 4}
Rounds 10
B.2 Hyperparameter Search Space for GLUE Benchmark
We perform hyperparameter (HP) tuning on RoBERTa-large separately for each GLUE task, selecting the optimal settings based on validation performance using a fixed random seed of 42. Final evaluations are conducted using two additional unseen seeds, {43, 44}, to ensure robustness. Table LABEL:tab:appendix_roberta_hyper depicts this.
Table 9: Hyperparameter(HP) settings of RoBERTa-large models on selected GLUE tasks for FedReFT, inspired from Wu et al. (2024b)
HP MNLI SST-2 QNLI QQP
position $p$ $p1$ $p3$ $p11$ $p11$
Tied weight False
Rank $r$ 1
Layer $L$ all
Dropout 0.05
Optimizer AdamW
LR $2\times 10^{-2}$
Weight decay 0.00
LR scheduler Linear
Batch size 32
Warmup ratio 0.00 0.10 0.10 0.06
Epochs 5
Rounds 50
Appendix C Dataset Description
C.1 Commonsense Reasoning
We train and evaluate our models on eight commonsense reasoning datasets spanning different types of open-ended QA tasks, following Hu et al. (2021a), we construct all examples. Table 10 shows the dataset samples.
  • BoolQ Clark et al. (2019): A yes/no question answering dataset consisting of naturally occurring questions. We remove the associated passages to ensure a fair comparison.
  • PIQA Bisk et al. (2020): A dataset for physical commonsense reasoning. The model must select the more plausible solution to everyday physical tasks.
  • SIQA Sap et al. (2019): Focuses on social interaction reasoning by asking the model to choose responses based on human intent and consequences.
  • HellaSwag Zellers et al. (2019): Requires choosing the most coherent sentence completion given a context, often involving physical or temporal common sense.
  • WinoGrande Sakaguchi et al. (2021): Inspired by the Winograd Schema Challenge Levesque et al. (2012), this dataset contains fill-in-the-blank problems with binary choices requiring commonsense coreference reasoning.
We follow the experimental setup in Hu et al. (2021a) by fine-tuning our models on a combined training corpus referred to as COMMONSENSE170K, which merges all of the above datasets. Evaluation is conducted individually on each dataset’s test split.
Table 10: Examples from commonsense reasoning tasks: BoolQClark et al. (2019), PIQABisk et al. (2020), HellaSwagZellers et al. (2019), and SIQASap et al. (2019). Each instruction is followed by the answer selected during evaluation.
Dataset Instruction / Question Answer
BoolQ Please answer the following question with true or false:
Question: Do Iran and Afghanistan speak the same language?
True
PIQA Please choose the correct solution to the question:
Question: When boiling butter, when it’s ready, you can
Solution1: Pour it onto a plate
Solution2: Pour it into a jar
Solution2
HellaSwag Please choose the correct ending to complete the given sentence:
Removing ice from car: Then, the man writes over the snow covering the window of a car, and a woman wearing winter clothes smiles. then
Ending1: , the man adds wax to the windshield and cuts it.
Ending2: , a person boards a ski lift…
Ending3: , the man puts on a christmas coat…
Ending4: , the man continues removing the snow on his car.
Ending4
SIQA Please choose the correct answer to the question:
Cameron decided to have a barbecue and gathered her friends together.
How would others feel as a result?
Answer1: like attending
Answer2: like staying home
Answer3: a good friend to have
Answer1
Table 11: Examples from math reasoning tasks: AQUA Ling et al. (2017), GSM8K Cobbe et al. (2021), and SVAMP Patel et al. (2021). Each instruction is followed by the correct answer derived through step-by-step reasoning.
Dataset Instruction / Question Answer
AQUA Solve the following word problem:
A car is driven in a straight line toward the base of a vertical tower. It takes 10 minutes for the angle of elevation to change from 45° to 60°. After how much more time will the car reach the base of the tower?
Answer Choices: (A) 5($\sqrt{3}$ + 1), (B) 6($\sqrt{3}$ + $\sqrt{2}$), (C) 7($\sqrt{3}$ – 1), (D) 8($\sqrt{3}$ – 2), (E) None of these.
GSM8K Solve the following question:
Janet’s ducks lay 16 eggs per day. She eats 3 eggs and uses 4 for baking. She sells the rest at \$2 per egg. How much money does she make daily?
\$18
SVAMP Solve the following arithmetic question:
Each pack of DVDs costs \$76. A discount of \$25 is applied. What is the final price per pack?
\$51
C.2 Arithmetic Reasoning
We evaluate arithmetic reasoning using seven benchmark datasets that cover a range of math word problem types. As in Hu et al. (2021a), we construct all examples without using golden or retrieved passages. Data samples are shows in Table 11.
  • AQuA Ling et al. (2017): Presents algebraic word problems in a multiple-choice format.
  • GSM8K Cobbe et al. (2021): A widely used benchmark of grade-school math problems requiring multi-step reasoning.
  • SVAMP Patel et al. (2021): A more challenging dataset that tests robustness to paraphrased and structurally altered word problems.
Following Hu et al. (2021a), we train our models on a combined training set named MATH10K.
C.3 Natural Language Understanding
For NLU, we evaluate on the GLUE benchmark following the evaluation protocol in Wu et al. (2024b). Data samples for shown in Table 12.
  • The validation set is split into two subsets one for in-training evaluation and the other for final testing.
  • For large datasets (QQP, MNLI, QNLI), 1,000 samples are used for in-training validation.
  • For smaller datasets, half of the validation set is used during training.
Table 12: Examples from GLUE benchmark  Wang et al. (2018) tasks: MNLI, SST-2, QNLI, and QQP.
Dataset Instruction / Question Answer
MNLI Premise: The dog is running through the field.
Hypothesis: An animal is moving.
Label: entailment
Entailment
SST-2 Sentence: A touching and thought-provoking piece of cinema.
Label: positive
Positive
QNLI Question: What is the capital of France?
Sentence: Paris is the capital and most populous city of France.
Label: entailment
Entailment
QQP Question1: How do I learn to play guitar?
Question2: What is the best way to learn guitar?
Label: duplicate
Duplicate
Appendix D Evaluation Loss-Based $\alpha$ Tuning
To personalize the blending of local and aggregated intervention parameters in FedReFT, we employ an evaluation loss-based tuning strategy for the mixing coefficient $\alpha\in[0,1]$. This coefficient governs how much each client integrates the aggregated All-But-Me (ABM) intervention parameters with its own local updates:
$\theta_{k}^{\text{new}}=(1-\alpha)\cdot\theta_{k}^{\text{local}}+\alpha\cdot\theta_{k}^{\text{ABM}},$ (10)
where $\theta_{k}$ can represent the LoReFT intervention components $W$, $R$, and $b$ for client $k$.
Tuning Procedure:
  1. 1.
    For a set of candidate $\alpha$ values ($\alpha\in\{0.0,0.1,\ldots,1.0\}$), the client computes interpolated parameters.
  2. 2.
    For each $\alpha$, the client evaluates its performance using a local validation set and records the evaluation loss $\mathcal{L}(\alpha)$.
  3. 3.
    The optimal mixing coefficient is selected by:
    $\alpha^{*}=\arg\min_{\alpha}\mathcal{L}(\alpha).$ (11)
Appendix E Theoretical Foundation: Geometric Median via Weiszfeld’s Algorithm
The geometric median offers a robust alternative to the arithmetic mean, particularly suitable for federated settings with heterogeneous or noisy client updates. For a given set of vectors $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}\subset\mathbb{R}^{d}$, the geometric median $\mathbf{y}^{*}$ is defined as:
$\mathbf{y}^{*}=\arg\min_{\mathbf{y}\in\mathbb{R}^{d}}\sum_{i=1}^{n}\|\mathbf{y}-\mathbf{x}_{i}\|_{2}.$ (12)
This optimization is non-smooth and convex, and generally lacks a closed-form solution. However, Weiszfeld’s algorithm Weiszfeld (1937) provides an efficient iterative method to approximate $\mathbf{y}^{*}$. We now derive and justify this algorithm via the Majorization-Minimization (MM) framework.
We define the cost function to be minimized:
$f(\mathbf{y})=\sum_{i=1}^{n}\|\mathbf{y}-\mathbf{x}_{i}\|_{2}.$ (13)
This function is convex but non-differentiable at points where $\mathbf{y}=\mathbf{x}_{i}$. Weiszfeld’s algorithm avoids such points during updates by construction.
The MM algorithm minimizes a difficult objective $f(\mathbf{y})$ by iteratively minimizing a surrogate function $Q(\mathbf{y}|\mathbf{y}^{(k)})$ that: Majorizes $f$: $Q(\mathbf{y}|\mathbf{y}^{(k)})\geq f(\mathbf{y})$ for all $\mathbf{y}$, Touches $f$ at the current iterate: $Q(\mathbf{y}^{(k)}|\mathbf{y}^{(k)})=f(\mathbf{y}^{(k)})$.
We define the surrogate using Jensen’s inequality and the convexity of the norm:
$Q(\mathbf{y}|\mathbf{y}^{(k)})=\sum_{i=1}^{n}\frac{\|\mathbf{y}-\mathbf{x}_{i}\|_{2}^{2}}{2\|\mathbf{y}^{(k)}-\mathbf{x}_{i}\|_{2}}+C(\mathbf{y}^{(k)}),$ (14)
where $C(\mathbf{y}^{(k)})$ is a constant that does not depend on $\mathbf{y}$. This function is differentiable and strictly convex in $\mathbf{y}$.
To find the minimizer of $Q(\mathbf{y}|\mathbf{y}^{(k)})$, we take the gradient and set it to zero:
$\nabla Q(\mathbf{y})=\sum_{i=1}^{n}\frac{\mathbf{y}-\mathbf{x}_{i}}{\|\mathbf{y}^{(k)}-\mathbf{x}_{i}\|_{2}}=0.$ (15)
Solving the above yields the Weiszfeld update rule:
$\mathbf{y}^{(k+1)}=\frac{\sum_{i=1}^{n}\frac{\mathbf{x}_{i}}{\|\mathbf{y}^{(k)}-\mathbf{x}_{i}\|_{2}}}{\sum_{i=1}^{n}\frac{1}{\|\mathbf{y}^{(k)}-\mathbf{x}_{i}\|_{2}}}.$ (16)
The update is only valid when $\mathbf{y}^{(k)}\neq\mathbf{x}_{i}$ for all $i$, a condition that can be enforced by initialization and step-size dampening if needed.
From MM theory Lange (2016), each iteration satisfies:
$f(\mathbf{y}^{(k+1)})$ $\leq Q(\mathbf{y}^{(k+1)}|\mathbf{y}^{(k)})$ (17)
$\leq Q(\mathbf{y}^{(k)}|\mathbf{y}^{(k)})=f(\mathbf{y}^{(k)}),$
ensuring that $f(\mathbf{y}^{(k)})$ is non-increasing. Under mild conditions (excluding cases where $\mathbf{y}^{(k)}=\mathbf{x}_{i}$), Weiszfeld’s algorithm converges to the geometric median $\mathbf{y}^{*}$.
E.1 Application in FedReFT: ABM Aggregation
In our FedReFT framework, each client receives an All-But-Me (ABM)aggregated update for intervention parameters computed as the geometric median of the corresponding parameters from all other clients. For client $k$, the ABM aggregated parameter is:
$\mathbf{W}^{\text{ABM}}_{k}=\arg\min_{\mathbf{w}\in\mathbb{R}^{d}}\sum_{m\neq k}\|\mathbf{w}-\mathbf{W}_{m}^{\text{local}}\|_{2}.$ (18)
We compute this using Weiszfeld’s algorithm for each parameter type independently, ensuring robustness to outlier clients and misaligned updates. This enables stable and personalized aggregation without sacrificing task-specific semantics.
Weiszfeld’s algorithm provides a theoretically grounded and computationally efficient way to compute the geometric median, making it ideal for ABM aggregation in heterogeneous FL. By leveraging this algorithm in FedReFT, we ensure robustness in aggregation and improve both convergence and personalization in non-i.i.d. federated environments.
Appendix F Ablation Study
F.1 Comparison of Aggregation method on different task
We use only 20% of the training data from the COMMONSENSE170K dataset, split among three clients, and evaluate the models using the SIQA task. Figure 3 shows that Geometric Median ABM aggregation outperforms all other approaches. Similarly, we split 50% of the MATH10K dataset among five clients, train each client for only five local epochs, and evaluate the results using the GSM8K evaluation set. Additionally, we use RTK GLUE for the natural language understanding (NLU) task.
Figure 3: Comparison of Aggregation Strategies Across Tasks. Performance of FedAvg, Mean-ABM, and Geometric Median-ABM on three benchmark task groups: Commonsense Reasoning, Arithmetic Reasoning, and GLUE. Geometric Median-ABM consistently outperforms FedAvg and Mean-ABM, highlighting its robustness and effectiveness in heterogeneous federated settings.
Table 13: Trainable Intervention Parameters across Models (in Millions) in FedReFT
Model Total P(M) # TP(M) TP%
LLaMa-1.1B 1100.05 0.72 0.0655
LLaMA 7B 6,738.42 2.10 0.0311
LLaMA-2 7B 6,738.42 2.10 0.0311
LLaMA-3 8B 8,030.27 2.10 0.0261
LlaMa-2-13B 13,015.86 6.55 0.0503
RoBERTa Large 355.36 0.0492 0.0138
Table 14: We vary LLaMA model sizes with $C=3$ clients following the DT design for commonsense reasoning task, alongside a centralized LoReFT baseline. As model capacity increases, we observe notable performance gains, with the largest model approaching the accuracy of the centralized setting. First four experiments on the Standalone centralize setup and later four experiments on the FL setup.
Method BoolQ PIQA HellaS. Avg $\uparrow$
LLaMa 7b 69.30 84.4 93.1 82.27
LLaMa-2 7B 71.10 83.8 94.3 83.07
LLaMa-3 7B 75.1 90.2 96.3 87.20
Tiny LLaMa 1B 63.83 49.18 46.03 53.01
LLaMa 7B 65.84 77.75 67.64 70.41
LLaMa-2 7B 68.93 74.81 77.73 73.82
LLaMa-3.2 3B 64.12 77.47 81.40 74.33
LLaMa-3 7B 72.60 85.85 89.85 82.77
Appendix G Communication Efficiency
As shown in Table LABEL:tab:appendix_trainable_param, FedReFT is communication and computationally efficient as it uses only a very small percentage of trainable parameters (TP) compared to the total model parameters. For example, in LLaMA-7B and LLaMA-2 7B, only 0.0311% of the total parameters are trained. In RoBERTa Large, this number is even smaller, at just 0.0138%. Even for large models like LLaMA-2-13B, the trainable portion remains as low as 0.0503%. This shows that FedReFT is highly parameter-efficient. Despite using such a small fraction of parameters, FedReFT still achieves strong performance, as discussed in the experimental analysis section 4. This highlights the benefit of using FedReFT in resource-constrained or communication-limited federated learning settings.
G.1 Additional Experimental Validation
In this section, we also conducted some additional experiments to show the robustness of FedReFT in different setups. Appendix Table LABEL:tbl:second_setup_commonsense depicts these.